A man and his brother have a combined age of 41 years. 3 years ago, the man was 6 times as old as his brother. how old is the man? and how old is the brother?
Answer by mananth(16946)

let man's age be x years
the brother's age = 41-x years
..
3 years ago
their ages were
Man x-3 years
brother 41-x-3= 38-x
..
equation
x-3=6(38-x)
x-3=228-6x
7x=231
x=33 years --man's age
brother's age = 41-33=8 years